OpenStudy (anonymous):

who needs help for math (algebra 1 mainly)

OpenStudy (anonymous):

i will

OpenStudy (goformit100):

Welcome To Open Study......... Sir @austinisawesome12

OpenStudy (amistre64):

how do you define 'help' for a 2.04 quiz?

OpenStudy (anonymous):

ill send a pic

OpenStudy (anonymous):

trying to substitute x

OpenStudy (amistre64):

a screenshot will provide more useful than some .webarchive file extention

OpenStudy (anonymous):

how do u screen shot on a iMac

OpenStudy (amistre64):

lol, good question

OpenStudy (anonymous):

thats why i did that

OpenStudy (amistre64):

command shift 3

OpenStudy (anonymous):

k

OpenStudy (amistre64):

seems like command control shift 3 saves your screen to the cilpboard, which can then be pasted into an image file. but i think command shift 3 copies and saves all at once

OpenStudy (anonymous):

it didn't do it

OpenStudy (amistre64):

hmm, then the only other idea i have is a camera

OpenStudy (amistre64):

OR if you can manually replicate it, then you might just want to type it all out

OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation.

OpenStudy (anonymous):

search that question up on open study it pops up right away

OpenStudy (amistre64):

so this is what it looks like

OpenStudy (anonymous):

yes i was trying to do that

OpenStudy (amistre64):

what is the top part of that suppose to mean?

OpenStudy (anonymous):

i know I'm supposed to substitute x for 1

OpenStudy (amistre64):

the top few lines are nothing but gibberish without any sort of context as to what its trying to convey.

OpenStudy (anonymous):

all i know is they are all supposed to equall each other and when it says 28x-1 they forgot the times 4

OpenStudy (anonymous):

thats what I'm trying to figure out to

OpenStudy (amistre64):

are they spose to be fractions maybe? i see that they multiplied the left by 3 and the right by 4 assuming they want a common denominator of 24

OpenStudy (anonymous):

its 3x+6 over 8

OpenStudy (anonymous):

and the other one over 6

OpenStudy (anonymous):

before each of those there is a 24 over 1

OpenStudy (amistre64):

\[\frac{3x+6}{8}=\frac{7x-1}{6}\] \[\frac{3}{3}\frac{3x+6}{8}=\frac{4}{4}\frac{7x-1}{6}\] \[\frac{3(3x+6)}{24}=\frac{4(7x-1)}{24}\] these are only equal if the top parts are equal so 3(3x+6) = 4(7x-1)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how do u use a fraction bar

OpenStudy (anonymous):

close tho

OpenStudy (amistre64):

lol, that an even odder question. i use it to indicate division

OpenStudy (anonymous):

nothings over 24

OpenStudy (amistre64):

we can modify the stuff, by multiplying each side by a useful form of '1' 3/3 = 1 and 4/4 = 1 in both instances we can get a common denominator in order to evaluate the numerators

OpenStudy (anonymous):

there is no 3/3 or 4/4 take this away

OpenStudy (anonymous):

the 24 is over that 1

OpenStudy (amistre64):

then he simply multipled both sides by 24/1

OpenStudy (anonymous):

24 (3x+6)= 24 (7x−16) 1 8 1 6

OpenStudy (amistre64):

either way we get the same results: \[\frac{3x+6}{8}=\frac{7x-1}{6}\] \[\frac{24}1\frac{3x+6}{8}=\frac{24}1\frac{7x-1}{6}\]

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

and then the rest of the problem

OpenStudy (amistre64):

in ascii, you can simply use the - key to represent a fraction 24 (3x+6) 24 (7x−16) -- ----- = -- ------- 1 8 1 6

OpenStudy (anonymous):

im 13 and doing algebra 1 its getting hard on me in 8th grade I'm doing geometry

OpenStudy (amistre64):

you agree that 24/8 = 3 and 24/6 = 4 ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so all that top part, as confusing as it is ... is fine. i see a mistake after it tho, do you see it as well?

OpenStudy (anonymous):

well i know that = 28x - 1 they forgot the times 4

OpenStudy (anonymous):

they didn't times 4 by 1

OpenStudy (amistre64):

right, they distributed correctly on the left side, but did not do it correctly on the right side. c(x+y) = cx + cy NOT c(x+y) = cx + y

OpenStudy (anonymous):

yes i got than

OpenStudy (anonymous):

that*

OpenStudy (amistre64):

but the question asks if the answer he got was correct. how do we determine the correctness of a result?

OpenStudy (anonymous):

replace x with 1

OpenStudy (amistre64):

yep, and after that, you might have learned something called cross multiplication. \[\frac{a}{b}=\frac{x}{y}~\text{ if and only if }~ay=bx\]

OpenStudy (anonymous):

yes but how do u get x by it self

OpenStudy (amistre64):

my example was in no means meant to give any special meaning to the letter x. but if we are to correct his mistake, how would you proceed?

OpenStudy (anonymous):

well since i know that 1 isn't the correct anwer would i divide by each side

OpenStudy (amistre64):

lets start here: 3(3x + 6) = 4(7x - 1) you agree that we distribute thru the paranthesis right? 3(3x + 6) = 4(7x - 1) 3(3x) + 3(6) = 4(7x) - 4(1) 9x + 18 = 28x - 4 correct?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

now it would be helpful if one side had no x part at all in it. how do we remove something in math?

OpenStudy (anonymous):

well thats what I'm having trouble on i think u divide to cancel out

OpenStudy (amistre64):

division doesnt cancel anything out .... division lets us modify things, but it never zeros anything out. what operation allows us to zero something out, to remove it completly.

OpenStudy (amistre64):

if i have a pie, and divide it into 1000 pieces, i am still left with some pie. division doesnt remove something.

OpenStudy (anonymous):

minus i am confused cause if u multiply it would just get bigger.......

OpenStudy (amistre64):

minus is correct. subtraction is a proper name for the operation :) lets remove 9x from the left side .... but in order to keep things equal, we would hav eto remove 9x from the right side as well, correct?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

9x + 18 = 28x - 4 -9x -9x ------------------ 18 = 19x - 4 now that we have one side clean of all its x parts, we want to deal with the other side. do you agree that the -4 is getting in the way of a solution? how would we remove it?

OpenStudy (anonymous):

you take -4 away from both sides of the equation

OpenStudy (amistre64):

correct. --4 ... which is just another way to say +4 since -4+4 = 0 18 = 19x - 4 +4 + 4 ------------ 22 = 19x now we are left with 19 mulitples of x on one side, and we dont want to remove the xs, we simply want to know what 1x is equal to. now we can divide.

OpenStudy (anonymous):

divide?

OpenStudy (amistre64):

yes, divide. you do know what division is right?

OpenStudy (anonymous):

lol yes just wondering what is there to divide

OpenStudy (amistre64):

well, if we have 19x .... but only want 1x; then 19x/by something, is equal to 1x what property do we know that allows us to make 19/? equal to 1

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

if we agree that 19 divides 19 equally into parts of 1; then 19x/19 = 1x

OpenStudy (anonymous):

yes thats right

OpenStudy (amistre64):

if we divide one side by 19, we have to divide the other side by 19 as well so ... 22 = 19x /19 /19 ---------- 22/19 = x

OpenStudy (anonymous):

where did the 22 come from

OpenStudy (amistre64):

you are spose to be following along ... and if anything, its the results of the last step we took :/ 18 = 19x - 4 +4 + 4 ------------ 22 = 19x

OpenStudy (anonymous):

oh ok yes i get it now sorry just got lost

OpenStudy (amistre64):

:) write it all out so you can view it as a whole process :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so, now we know what an x is equal to: 22/19

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so thats the answer

OpenStudy (amistre64):

that is one form of the answer. but it is accurate enough. i see now way of reducing it, and im not sure of a mixed numeral is what they are looking for.

OpenStudy (anonymous):

its right thats what i got the first time ii ever did this but then i thought i was wayyy off

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=3%283x%2B6%29%3D4%287x-1%29 the wolf likes it :)

OpenStudy (anonymous):

how do i give u a medal

OpenStudy (anonymous):

and why do i have a 12 next to my name

OpenStudy (amistre64):

if you do not see a 'best response' button next to all of my posts, then you might need to refresh your browser to see it. and im assuming you have a 12 next to your name becuase that is the username you choose.

OpenStudy (anonymous):

there gave ya a medal

OpenStudy (amistre64):

'mathematics' and 'ashleyhart' are the only other accounts associated with your IP address.

OpenStudy (anonymous):

now I'm 13

OpenStudy (amistre64):

thnx for the medal :) well, im not 64 yet .... give me a few years lol